Create a real-life situation that fits into the equation (x+4)(x-7) = 0 and express the situation as the same equation.
Answer by ankor@dixie-net.com(22740)

Create a real-life situation that fits into the equation (x+4)(x-7) = 0 and express the situation as the same equation.
:
A simple problems would be. You want a rectangular area of 28 sq/ft
You want the length to be 3 ft longer than the width, Find the dimensions:
;
Let x = length
then
(x-3) = width
and
Area = x(x-3) = x^2 - 3x
:
Area equation:
x^2 - 3x = 28
:
A quadratic equation
x^2 -3x - 28 = 0
:
factor this to:
(x-7)(x+4) = 0
:
Positive solution:
x = +7 is the length